HELIUM NANODROPLET ISOLATION OF IONIC LIQUID VAPOR: INRARED LASER SPECTROSCOPY OF [EMIM][Tf_{2}N]

\underline{STEVEN D. FLYNN}, GARY E.DOUBERLY, Department of Chemistry, University of Georgia, Athens, Georgia, USA 30602.

The Infrared spectrum of the vapor produced upon thermal vaporization of the [emim][Tf_{2}N]^{a} ionic liquid has been obtained using the helium nanodroplet isolation method. Despite the low vapor pressure of [emim][Tf_{2}N], sufficient gas phase densities are produced, allowing for efficient helium nanodroplet pick-up. The mass spectrum of the emim[Tf_{2}N] doped droplet beam shows signatures^{b} that have been attributed in gas phase measurements to the presence of isolated, intact [emim][Tf_{2}N] ion-pairs. Furthermore, the mass spectrometry results indicate that emim[Tf_{2}N] does not undergo thermal decomposition at 410 K. Comparisons are made between the experimental measurements and ab initio calculations (mp2/6-311++g(d,p)) of the CH stretch vibrational bands and permanent electric dipole moments for several [emim][Tf_{2}N] low energy isomers. The helium nanodroplet infrared spectrum of this species provides rather definitive support to the previously suggested vaporization mechanism of ionic liquids.^{c}

a[emim][Tf_{2}N] is defined as 1-ethyl-3-methylimidazolium[bis(trifluoromethylsulfonyl)imide]

bArmstrong, J.P.; Hurst, C.; Jones, R. G.; Licence, P.; Lovelock, K. R. J.; Satterley, C. J.; Villar-Garcia, I. J. Physical Chemistry Chemical Physics 2007, 9, 982.

Strasser, D.; Goulay, F.; Belau, L.; Kostko, O.; Koh, C.; Chambreau, S. D.; Vaghjiani, G. L.; Ahmed, M.; Leone, S. R. textitJournal of Physical Chemistry A 2010, textit114, 879.

cStrasser, D.; Goulay, F.; Kelkar, M. S.; Maginn, E. J.; Leone, S. R. Journal of Physical Chemistry A 2007, 111, 3191.

Chambreau, S. D.; Vaghjiani, G. L.; To, A.; Koh, C.; Strasser, D.; Kostko, O.; Leone, S. R. Journal of Physical Chemistry B 2010, 114, 1361.

Maginn, E. J.; Kelkar, M. S. Journal of Physical Chemistry B 2007, 111, 9424.